# Energy landscape analysis based on the Ising model: Tutorial review

### v1

@@ 1 Introduction
Energy landscape is an established concept, catered with computational methods, in protein folding research [1, 2, 3, 4]. With this dynamic view, we comprehend protein folding as an ensemble of locally stable protein conformations that are stochastically switched between depending on how stable each conformation is and how likely one conformation changes to another. According to the principles of statistical physics, a conformation with a low energy occurs with a high probability, and transitions between conformations occur with a high frequency if the energy barrier separating them is low (see Figs. 1 and 2 for schematics). The purpose of this review paper is to explain a family of data analysis methods for constructing such energy landscapes and dynamics on them for general multivariate time series data for which we do not know any underlying kinetic, chemical, or other laws.
@@
Figure 1: One-dimensional schematic of energy landscapes. (a) An energy landscape with two local minima of energy, $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$. They are separated by an energy barrier. The filled circle, or “ball”, represents the state of the system, which tends to go downhill but can also go uphill with a smaller rate. The position of the ball along the horizontal axis specifies the system’s state, which we call the activity pattern. (b) An energy landscape with a higher energy barrier between two local minima. (c) Sample time course of the system’s dynamics on the energy landscape shown in (a). (d) Sample time course on the energy landscape shown in (b). Note that switching between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ is rarer in (d) than (c) because $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ are separated by a higher energy barrier in (b) than (a). In fact, the system’s state (i.e., filled circle) can take activity patterns different from the two local minima (e.g., a pattern close to but different from $\bm{\alpha}_{1}$). However, we only show stochastic transition dynamics between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ in (c) and (d) for simplicity.
@@
Figure 2: Two-dimensional schematic of an energy landscape. Four local minima are present. The dotted lines show the borders between the basin of different local minima. The “ball” represents the state of the system, and its position on the plane (ignoring the vertical coordinate) represents the activity pattern.
@@
The type of energy landscape analysis (ELA) that this article deals with, which we simply refer to as ELA in the following text, was initially proposed in [5, 6] for functional magnetic resonance imaging (fMRI) data recorded from the human brain. It uses the Ising spin system model. The model fitting part of this ELA is shared with earlier studies, in particular in neuroscience [7, 8, 9, 10, 11, 12, 13]. Related to this fact, the ELA has predominantly been used in neuroscience, especially for human fMRI data [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Other applications in neuroscience include electroencephalogram (EEG) [36, 37, 38], magnetoencephalography (MEG) [39], and structural brain network [40] data. However, the ELA does not impose any neuroscientific assumptions on the data and therefore applicable to any multivariate time series data as long as the data are long enough relative to the number of variables. In fact, the ELA has seen applications to simulated data of cardiac fibrillation [41], time series of a combination of multimodal observables from patients in rheumatoid arthritis (such as the rheumatoid factor and swollen joint count) [42], Raman spectroscopy data [43], microbiome data [44, 45, 46, 47], and mixture of microbiome data and physiological signals from calves [48]. ELA crucially takes a system’s dynamics point of view in the sense that the method tracks the dynamics of the system’s state, which is defined by the values of all variables at each time $t$. Such a system’s state is abstracted into the filled ball in Figs. 1(a), 1(b), and 2. As these figures show, ELA aims to capture the given multivariate time series data as the single abstracted system’s state that transits among discrete states, similar to a Markov chain. Furthermore, the present ELA is based on the computation and statistical physics theory of the Ising spin system model. There are various other energy landscape analyses in addition to those in protein folding studies (e.g. [49, 50, 51, 52, 53]). In this review paper, we exclusively focus on the ELA based on the Ising model.
@@ 2.1 Overview
We outline a workflow of ELA in Fig. 3. First, the input data are multivaraite time series with $N$ variables in discrete time (Fig. 3(a)). We denote the time series by $\{\bm{x}(1),\ldots,\bm{x}(t_{\max})\}$, where $\bm{x}(t)$ (with $t\in\{1,\ldots,t_{\max}\}$) is the $N$-dimensional time series at time $t$, and $t_{\max}$ is the number of observations, or the length of the time series. We denote the $N$-dimensional vector $\bm{x}(t)$ by
@@
Figure 3: Steps of the ELA. (a) The input data is multivariate time series with $N$ variables. (b) Binarized time series. (c) Fraction of times, $P(\bm{\sigma})$, with which each of the $2^{N}$ activity patterns, $\bm{\sigma}$, occurs in the binarized data shown in (b). (d) Fitting of the PMEM to the distribution of $\bm{\sigma}$ shown in (c). (e) Disconnectivity graph derived from the estimated PMEM. It shows relationships between the activity patterns that are local minima. Each $\bm{\alpha}_{i}$, $i\in\{1,\ldots,6\}$, represents a local minimum. The double-headed arrow indicates the height of the energy barrier between local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ when the activity pattern tries to travel from $\bm{\alpha}_{1}$ to $\bm{\alpha}_{2}$. (f) Basin graph derived from the estimated PMEM. (g) Downstream analysis, which may use a feature computed from the estimated disconnectivity graph or basin graph.
Second, we binarize each $x_{i}(t)$, $i\in\{1,\ldots,N\}$, $t\in\{1,\ldots,t_{\max}\}$. See Fig. 3(b) for an example of binarized multivariate time series. We denote the obtained binarized multivariate time series by $\{\bm{\sigma}(1),\ldots,\bm{\sigma}(t_{\max})\}$, where the system’s binarized state, which we call the activity pattern, at time $t$ is written as
@@
where $\sigma_{i}(t)\in\{-1,1\}$. Another common convention is $\sigma_{i}(t)\in\{0,1\}$. In either convention, $1$ corresponds to high activity of the $i$th variable at time $t$; the other value (i.e., $-1$ or 00 depending on the coding scheme) corresponds to low activity. See [54] for discussion of the two conventions.
Third, we estimate the Ising model, i.e., infer the model parameters from the binarized multivariate time series data $\{\bm{\sigma}(1),\ldots,\bm{\sigma}(t_{\max})\}$ (Fig. 3(c) and (d)). The Ising model is also referred to the pairwise maximum entropy model (PMEM), which we use throughout this paper, and the Boltzmann machine. There are various inference algorithms. The estimated Ising model is typically a maximum likelihood estimator for the data under some constraints. In fact, the procedures up to this point, i.e., fitting the Ising model to multivariate time series and other multivariate data, is not a new idea. See [11, 12, 55, 56, 57, 58] for reviews.
The following steps are unique to the ELA. Fourth, we estimate the so-called disconnectivity graph [59], of which a hypothetical example is shown in Fig. 3(e). The bottom of each valley schematically shown in Figs. 1 and 2 corresponds to the tip of each branch hanging down vertically in the disconnectivity graph (see Fig. 3(e)). It is called a local minimum of energy (local minimum for short). The disconnectivity graph represents all the local minima and their relationships in terms of the ease of transition for the estimated PMEM and is a primary product of the ELA. Another main product of the ELA is the basin graph, shown in Fig. 3(f). The basin graph encodes the width or area of the valley of each local minimum schematically shown in Figs. 1 and 2 and contains other information.
@@
Sixth, one often wants to compare the energy landscapes obtained from a group of people or an individual against the energy landscapes obtained from a different group or individual. Such a comparison is made in terms of features of the energy landscape. For example, we examine whether or not larger values of a particular feature from the energy landscape estimated for different human individuals are associated with the severity of a disease (see Fig. 3(g)). In the following sections, we explain each step in detail.
@@ 2.2 Fitting the PMEM
We start by binarizing the original time series (see Fig. 3(a) and (b)). For each $i\in\{1,\ldots,N\}$ and $t\in\{1,\ldots,t_{\max}\}$, we set $\sigma_{i}(t)=-1$ if $x_{i}(t)<\theta_{i}$ and $\sigma_{i}(t)=1$ if $x_{i}(t)\geq\theta_{i}$, where $\theta_{i}$ is an arbitrary threshold. A common choice of $\theta_{i}$ is the 50 percentile of $\{x_{i}(1),\ldots,x_{i}(t_{\max})\}$ such that $\sigma_{i}(t)=-1$ and $\sigma_{i}(t)=1$ with probability $0.5$ each. If the distribution of $x_{i}(t)$ is not extremely skewed, one can also set $\theta_{i}$ to the average of the time series, i.e., $\sum_{t=1}^{t_{\max}}x_{i}(t)/t_{\max}$, to obtain roughly 50% of $\sigma_{i}(t)=-1$ and $\sigma_{i}(t)=1$ each.
Because $\bm{\sigma}(t)$ is an $N$-dimensional binary vector, there are $2^{N}$ possible activity patterns in total, which we enumerate as $\bm{\sigma}^{(1)}$, $\ldots$, $\bm{\sigma}^{(2^{N})}$. This is depicted in Fig. 3(c). For example, one obtains $\bm{\sigma}^{(1)}=(-1,-1,\ldots,-1)$, $\bm{\sigma}^{(2)}=(1,-1,\ldots,-1)$, and $\bm{\sigma}^{(2^{N})}=(1,1,\ldots,1)$. We then count the fraction of each activity pattern $\bm{\sigma}$, denoted by $P_{\text{data}}(\bm{\sigma})$, in the given data. Note that $\bm{\sigma}$ is any of $\bm{\sigma}^{(1)}$, $\ldots$, $\bm{\sigma}^{(2^{N})}$ and that $P_{\text{data}}(\bm{\sigma})$ is simply the number of times that $\bm{\sigma}$ appears in the data that is divided by $t_{\max}$. Because $P_{\text{data}}(\bm{\sigma})$ is a fraction, we obtain
@@
The PMEM aims to fit a particular form of the distribution, denoted by $P_{\text{model}}(\bm{\sigma})$ and explained in the following text, to the empirical distribution, $P_{\text{data}}(\bm{\sigma})$. With this methodological choice, we are drastically ignoring the temporal order of the given data. For example, the estimated PMEM will be the same as that for the original time series even if we re-arrange the original multivariate time series in the reversed order (i.e., $\{\bm{\sigma}(t_{\max}),\bm{\sigma}(t_{\max}-1),\ldots,\bm{\sigma}(1)\}$) or shuffle the time stamp of $\{\bm{\sigma}(t)\}$ randomly before the analysis. Therefore, the ELA is in a sense not a method for analyzing time series. It is a method to describe the distribution of the given data. However, ELA still allows us to infer and interpret the dynamics of the data; e.g., the system’s state (i.e., the ball in Figs. 1 and 2) stays near a local minimum $\bm{\alpha}_{1}$ for some time, then transits to a different local minimum $\bm{\alpha}_{2}$, and comes back to $\bm{\alpha}_{1}$. This is because the PMEM implies a particular equilibrium stochastic dynamics. In other words, once a PMEM is estimated, it dictates the probability or rate of the move from any activity pattern $\bm{\sigma}$ to another, which we explain in section 3.2.3.
@@
The PMEM maximizes the entropy of the distribution of activity patterns, $P(\bm{\sigma})$, under the condition that $\langle\sigma_{i}\rangle$ and $\langle\sigma_{i}\sigma_{j}\rangle$ (with $1\leq i\leq j\leq N)$ are the same between the estimated Ising model and the given data for all $i,j\in\{1,\ldots,N\}$. As a result of entropy maximization, one obtains (see Fig. 3(d))
@@
Equation (6) is a Boltzmann distribution (also called Gibbs distribution) in statistical mechanics without the temperature parameter, or equivalently, the temperature normalized to $1$. Equation (6) implies that a high probability of the activity pattern $P(\bm{\sigma})$ corresponds to a low energy $E(\bm{\sigma})$ and vice versa. It should be noted that the energy here is a purely mathematical construct and does not imply any physical energy. In Eq. (7), the fitting parameter $h_{i}$ represents the tendency that the $i$th variable is active (i.e., $\sigma_{i}=1$), and $J_{ij}$ quantifies the sign and strength of the pairwise interaction between the $i$th and $j$th variables. The PMEM assumes that $J_{ij}=J_{ji}$. Furthermore, the $J_{ij}\sigma_{i}\sigma_{j}$ when $i=j$ is reduced to a constant $J_{ii}\sigma_{i}^{2}=J_{ii}$ because $\sigma_{i}^{2}$ with $\sigma_{i}\in\{-1,1\}$ is always equal to $1$. Because $P(\bm{\sigma})$ in Eq. (6) does not change by the addition of a constant term to $E(\bm{\sigma})$, one can simplify Eq. (7) to
@@
where superscripts new and old represent the values after and before a single updating step, respectively, and $\epsilon$ is the learning rate. The factor $1/t_{\max}$ on the right-hand sides of Eqs. (11) and (12) is for normalization. A slightly different iteration scheme called the iterative scaling algorithm [60], in which the right-hand side of Eqs. (11) and (12) is replaced by $\epsilon\mathop{\mathrm{sgn}}(\langle\sigma_{i}\rangle)\log(\langle\sigma_{i}\rangle_{\text{data}}/\langle\sigma_{i}\rangle_{\rm{model}})$ and $\epsilon\mathop{\mathrm{sgn}}(\langle\sigma_{i}\sigma_{j}\rangle)\log(\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}/\langle\sigma_{i}\sigma_{j}\rangle_{\rm{model}})$, respectively, is also common [7, 11, 13]. With either iterative algorithm, the convergence (i.e., $h_{i}^{\text{new}}=h_{i}^{\text{old}}$ and $J_{ij}^{\text{new}}=J_{ij}^{\text{old}}$ for all $i,j$) implies that $\langle\sigma_{i}\rangle_{\text{model}}=\langle\sigma_{i}\rangle_{\rm data}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}=\langle\sigma_{i}\sigma_{j}\rangle_{\rm data}$ for all $i,j$.
One can verify that Eq. (6) is concave in terms of each $h_{i}$ and $J_{ij}$ by checking that the Hessian of $\log\mathcal{L}$ is a type of sign-flipped covariance matrix, which is negative semi-definite. The concavity guarantees that the gradient ascent scheme always yields the unique maximum likelihood estimator. However, a single updating step requires reference to $P(\bm{\sigma})$ for all the $2^{N}$ activity patterns for calculating $\langle\sigma_{i}\rangle_{\rm{model}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\rm{model}}$. Because $P(\bm{\sigma})$ changes in every iteration step, the gradient ascent method is computationally costly for large $N$. Our experience is that, without dedicated hardware, the gradient ascent method would be slow even with $N=15$ or 20. To circumvent this problem, there are various approximate algorithms (e.g., pseudo-likelihood maximization) for estimating the PMEM [61, 62, 55, 54, 56].

### v2

@@ 1 Introduction
Energy landscape is an established concept, catered with computational methods, in protein folding research [1, 2, 3, 4]. With this dynamic view, we comprehend protein folding as an ensemble of locally stable protein conformations that are stochastically switched between depending on how stable each conformation is and how likely one conformation changes to another. According to the principles of statistical physics, a conformation with a low energy occurs with a high probability, and transitions between conformations occur with a high frequency if the energy barrier separating them is low (see Figs 1 and 2 for schematics). The purpose of this review paper is to explain a family of data analysis methods for constructing such energy landscapes and dynamics on them for general multivariate time series data for which we do not know any underlying kinetic, chemical, or other laws.
@@
Fig 1: One-dimensional schematic of energy landscapes. (a) An energy landscape with two local minima of energy, $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$. They are separated by an energy barrier. The filled circle, or “ball”, represents the state of the system, which tends to go downhill but can also go uphill with a smaller rate. The position of the ball along the horizontal axis specifies the system’s state, which we call the activity pattern. (b) An energy landscape with a higher energy barrier between two local minima. (c) Sample time course of the system’s dynamics on the energy landscape shown in (a). (d) Sample time course on the energy landscape shown in (b). Note that switching between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ is rarer in (d) than (c) because $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ are separated by a higher energy barrier in (b) than (a). In fact, the system’s state (i.e., filled circle) can take activity patterns different from the two local minima (e.g., a pattern close to but different from $\bm{\alpha}_{1}$). However, we only show stochastic transition dynamics between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ in (c) and (d) for simplicity.
@@
Fig 2: Two-dimensional schematic of an energy landscape. Four local minima are present. The dotted lines show the borders between the basin of different local minima. The “ball” represents the state of the system, and its position on the plane (ignoring the vertical coordinate) represents the activity pattern.
@@
The type of energy landscape analysis (ELA) that this article deals with, which we simply refer to as ELA in the following text, was initially proposed in [5, 6] for functional magnetic resonance imaging (fMRI) data recorded from the human brain. It uses the Ising spin system model. The model fitting part of this ELA is shared with earlier studies, in particular in neuroscience [7, 8, 9, 10, 11, 12, 13]. Related to this fact, the ELA has predominantly been used in neuroscience, especially for human fMRI data [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Other applications in neuroscience include electroencephalogram (EEG) [36, 37, 38], magnetoencephalography (MEG) [39], and structural brain network [40] data. However, the ELA does not impose any neuroscientific assumptions on the data and therefore applicable to any multivariate time series data as long as the data are long enough relative to the number of variables. In fact, the ELA has seen applications to simulated data of cardiac fibrillation [41], time series of a combination of multimodal observables from patients in rheumatoid arthritis (such as the rheumatoid factor and swollen joint count) [42], Raman spectroscopy data [43], microbiome data [44, 45, 46, 47], and mixture of microbiome data and physiological signals from calves [48]. ELA crucially takes a system’s dynamics point of view in the sense that the method tracks the dynamics of the system’s state, which is defined by the values of all variables at each time $t$. Such a system’s state is abstracted into the filled ball in Figs 1(a), 1(b), and 2. As these figures show, ELA aims to capture the given multivariate time series data as the single abstracted system’s state that transits among discrete states, similar to a Markov chain. Furthermore, the present ELA is based on the computation and statistical physics theory of the Ising spin system model. There are various other energy landscape analyses in addition to those in protein folding studies (e.g. [49, 50, 51, 52, 53]). In this review paper, we exclusively focus on the ELA based on the Ising model.
@@ 2.1 Overview
We outline a workflow of ELA in Fig 3. First, the input data are multivaraite time series with $N$ variables in discrete time (Fig 3(a)). We denote the time series by $\{\bm{x}(1),\ldots,\bm{x}(t_{\max})\}$, where $\bm{x}(t)$ (with $t\in\{1,\ldots,t_{\max}\}$) is the $N$-dimensional time series at time $t$, and $t_{\max}$ is the number of observations, or the length of the time series. We denote the $N$-dimensional vector $\bm{x}(t)$ by
@@
Fig 3: Steps of the ELA. (a) The input data is multivariate time series with $N$ variables. (b) Binarized time series. (c) Fraction of times, $P(\bm{\sigma})$, with which each of the $2^{N}$ activity patterns, $\bm{\sigma}$, occurs in the binarized data shown in (b). (d) Fitting of the PMEM to the distribution of $\bm{\sigma}$ shown in (c). (e) Disconnectivity graph derived from the estimated PMEM. It shows relationships between the activity patterns that are local minima. Each $\bm{\alpha}_{i}$, $i\in\{1,\ldots,6\}$, represents a local minimum. The double-headed arrow indicates the height of the energy barrier between local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ when the activity pattern tries to travel from $\bm{\alpha}_{1}$ to $\bm{\alpha}_{2}$. (f) Basin graph derived from the estimated PMEM. (g) Downstream analysis, which may use a feature computed from the estimated disconnectivity graph or basin graph.
Second, we binarize each $x_{i}(t)$, $i\in\{1,\ldots,N\}$, $t\in\{1,\ldots,t_{\max}\}$. See Fig 3(b) for an example of binarized multivariate time series. We denote the obtained binarized multivariate time series by $\{\bm{\sigma}(1),\ldots,\bm{\sigma}(t_{\max})\}$, where the system’s binarized state, which we call the activity pattern, at time $t$ is written as
@@
where $\sigma_{i}(t)\in\{-1,1\}$. Another common convention is $\sigma_{i}(t)\in\{0,1\}$. In either convention, $1$ corresponds to high activity of the $i$th variable at time $t$; the other value (i.e., $-1$ or 00 depending on the coding scheme) corresponds to low activity. Because the linear transformation $f(x)=(x+1)/2$ maps $\{-1,1\}$ to $\{0,1\}$, i.e., $f(-1)=0$ and $f(1)=1$, the choice of either convention does not essentially matter for the following analyses. See [54] for more discussion of the two conventions.
Third, we estimate the Ising model, i.e., infer the model parameters from the binarized multivariate time series data $\{\bm{\sigma}(1),\ldots,\bm{\sigma}(t_{\max})\}$ (Fig 3(c) and (d)). The Ising model is also referred to the pairwise maximum entropy model (PMEM), which we use throughout this paper, and the Boltzmann machine. There are various inference algorithms. The estimated Ising model is typically a maximum likelihood estimator for the data under some constraints. In fact, the procedures up to this point, i.e., fitting the Ising model to multivariate time series and other multivariate data, is not a new idea. See [11, 12, 55, 56, 57, 58] for reviews.
The following steps are unique to the ELA. Fourth, we estimate the so-called disconnectivity graph [59], of which a hypothetical example is shown in Fig 3(e). The bottom of each valley schematically shown in Figs 1 and 2 corresponds to the tip of each branch hanging down vertically in the disconnectivity graph (see Fig 3(e)). It is called a local minimum of energy (local minimum for short). The disconnectivity graph represents all the local minima and their relationships in terms of the ease of transition for the estimated PMEM and is a primary product of the ELA. Another main product of the ELA is the basin graph, shown in Fig 3(f). The basin graph encodes the width of the valley of each local minimum schematically shown in Fig 1 or its area schematically shown in Fig 2. It also contains other information.
@@
Sixth, one often wants to compare the energy landscapes obtained from a group of people or an individual against the energy landscapes obtained from a different group or individual. Such a comparison is made in terms of features of the energy landscape. For example, we examine whether or not larger values of a particular feature from the energy landscape estimated for different human individuals are associated with the severity of a disease (see Fig 3(g)). In the following sections, we explain each step in detail.
@@ 2.2 Fitting the PMEM
We start by binarizing the original time series (see Fig 3(a) and (b)). For each $i\in\{1,\ldots,N\}$ and $t\in\{1,\ldots,t_{\max}\}$, we set $\sigma_{i}(t)=-1$ if $x_{i}(t)<\theta_{i}$ and $\sigma_{i}(t)=1$ if $x_{i}(t)\geq\theta_{i}$, where $\theta_{i}$ is an arbitrary threshold. A common choice of $\theta_{i}$ is the 50 percentile of $\{x_{i}(1),\ldots,x_{i}(t_{\max})\}$ such that $\sigma_{i}(t)=-1$ and $\sigma_{i}(t)=1$ with probability $0.5$ each. If the distribution of $x_{i}(t)$ is not extremely skewed, one can also set $\theta_{i}$ to the average of the time series, i.e., $\sum_{t=1}^{t_{\max}}x_{i}(t)/t_{\max}$, to obtain roughly 50% of $\sigma_{i}(t)=-1$ and $\sigma_{i}(t)=1$ each.
Because $\bm{\sigma}(t)$ is an $N$-dimensional binary vector, there are $2^{N}$ possible activity patterns in total, which we enumerate as $\bm{\sigma}^{(1)}$, $\ldots$, $\bm{\sigma}^{(2^{N})}$. This is depicted in Fig 3(c). For example, one obtains $\bm{\sigma}^{(1)}=(-1,-1,\ldots,-1)$, $\bm{\sigma}^{(2)}=(1,-1,\ldots,-1)$, and $\bm{\sigma}^{(2^{N})}=(1,1,\ldots,1)$. We then count the fraction of each activity pattern $\bm{\sigma}$, denoted by $P_{\text{data}}(\bm{\sigma})$, in the given data. Note that $\bm{\sigma}$ is any of $\bm{\sigma}^{(1)}$, $\ldots$, $\bm{\sigma}^{(2^{N})}$ and that $P_{\text{data}}(\bm{\sigma})$ is simply the number of times that $\bm{\sigma}$ appears in the data that is divided by $t_{\max}$. Because $P_{\text{data}}(\bm{\sigma})$ is a fraction, we obtain
@@
The PMEM aims to fit a particular form of the distribution, denoted by $P_{\text{model}}(\bm{\sigma})$ and explained in the following text, to the empirical distribution, $P_{\text{data}}(\bm{\sigma})$. With this methodological choice, we are drastically ignoring the temporal order of the given data. For example, the estimated PMEM will be the same as that for the original time series even if we re-arrange the original multivariate time series in the reversed order (i.e., $\{\bm{\sigma}(t_{\max}),\bm{\sigma}(t_{\max}-1),\ldots,\bm{\sigma}(1)\}$) or shuffle the time stamp of $\{\bm{\sigma}(t)\}$ randomly before the analysis. Therefore, the ELA is in a sense not a method for analyzing time series. It is a method to describe the distribution of the given data. However, ELA still allows us to infer and interpret the dynamics of the data; e.g., the system’s state (i.e., the ball in Figs 1 and 2) stays near a local minimum $\bm{\alpha}_{1}$ for some time, then transits to a different local minimum $\bm{\alpha}_{2}$, and comes back to $\bm{\alpha}_{1}$. This is because the PMEM implies a particular equilibrium stochastic dynamics. In other words, once a PMEM is estimated, it dictates the probability or rate of the move from any activity pattern $\bm{\sigma}$ to another, which we explain in Section 3.2.3.
@@
The PMEM maximizes the entropy of the distribution of activity patterns, $P(\bm{\sigma})$, under the condition that $\langle\sigma_{i}\rangle$ and $\langle\sigma_{i}\sigma_{j}\rangle$ (with $1\leq i\leq j\leq N)$ are the same between the estimated Ising model and the given data for all $i,j\in\{1,\ldots,N\}$. As a result of entropy maximization, one obtains (see Fig 3(d))
@@
Equation (6) is a Boltzmann distribution (also called Gibbs distribution) in statistical mechanics without the temperature parameter, or equivalently, the temperature normalized to $1$. Equation (6) implies that a high probability of the activity pattern $P(\bm{\sigma})$ corresponds to a low energy $E(\bm{\sigma})$ and vice versa. It should be noted that the energy here is a purely mathematical construct and does not imply any physical energy. In Eq (7), the fitting parameter $h_{i}$ represents the tendency that the $i$th variable is active (i.e., $\sigma_{i}=1$), and $J_{ij}$ quantifies the sign and strength of the pairwise interaction between the $i$th and $j$th variables. The PMEM assumes that $J_{ij}=J_{ji}$. Furthermore, the $J_{ij}\sigma_{i}\sigma_{j}$ when $i=j$ is reduced to a constant $J_{ii}\sigma_{i}^{2}=J_{ii}$ because $\sigma_{i}^{2}$ with $\sigma_{i}\in\{-1,1\}$ is always equal to $1$. Because $P(\bm{\sigma})$ in Eq (6) does not change by the addition of a constant term to $E(\bm{\sigma})$, one can simplify Eq (7) to
@@
where superscripts new and old represent the values after and before a single updating step, respectively, and $\epsilon$ is the learning rate. The factor $1/t_{\max}$ on the right-hand sides of Eqs (11) and (12) is for normalization. A slightly different iteration scheme called the iterative scaling algorithm [60], in which the right-hand side of Eqs (11) and (12) is replaced by $\epsilon\mathop{\mathrm{sgn}}(\langle\sigma_{i}\rangle)\log(\langle\sigma_{i}\rangle_{\text{data}}/\langle\sigma_{i}\rangle_{\rm{model}})$ and $\epsilon\mathop{\mathrm{sgn}}(\langle\sigma_{i}\sigma_{j}\rangle)\log(\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}/\langle\sigma_{i}\sigma_{j}\rangle_{\rm{model}})$, respectively, is also common [7, 11, 13]. With either iterative algorithm, the convergence (i.e., $h_{i}^{\text{new}}=h_{i}^{\text{old}}$ and $J_{ij}^{\text{new}}=J_{ij}^{\text{old}}$ for all $i,j$) implies that $\langle\sigma_{i}\rangle_{\text{model}}=\langle\sigma_{i}\rangle_{\rm data}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}=\langle\sigma_{i}\sigma_{j}\rangle_{\rm data}$ for all $i,j$.
One can verify that Eq (6) is concave in terms of each $h_{i}$ and $J_{ij}$ by checking that the Hessian of $\log\mathcal{L}$ is a type of sign-flipped covariance matrix, which is negative semi-definite. The concavity guarantees that the gradient ascent scheme always yields the unique maximum likelihood estimator. However, a single updating step requires reference to $P(\bm{\sigma})$ for all the $2^{N}$ activity patterns for calculating $\langle\sigma_{i}\rangle_{\rm{model}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\rm{model}}$. Because $P(\bm{\sigma})$ changes in every iteration step, the gradient ascent method is computationally costly for large $N$. Our experience is that, without dedicated hardware, the gradient ascent method would be slow even with $N=15$ or 20. To circumvent this problem, there are various approximate algorithms (e.g., pseudo-likelihood maximization) for estimating the PMEM [61, 62, 55, 54, 56].